# CultureGuard: Towards Culturally-Aware Dataset and Guard Model for Multilingual Safety Applications

### v2

@@ Abstract
The increasing use of Large Language Models (LLMs) in agentic applications highlights the need for robust safety guard models. While content safety in English is well-studied, non-English languages lack similar advancements due to the high cost of collecting culturally aligned labeled datasets. We present CultureGuard, a novel solution for curating culturally aligned, high-quality safety datasets across multiple languages. Our approach introduces a four-stage synthetic data generation and filtering pipeline: cultural data segregation, cultural data adaptation, machine translation, and quality filtering. This pipeline enables the conversion and expansion of the Nemotron-Content-Safety-Dataset-V2 Ghosh et al. (2025) English safety dataset into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese. The resulting dataset, Nemotron-Content-Safety-Dataset-Multilingual-v1, comprises 386,661 samples in 9 languages and facilitates the training of Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 via LoRA-based fine-tuning. The final model achieves state-of-the-art performance on several multilingual content safety benchmarks. We also benchmark the latest open LLMs on multilingual safety and observe that these LLMs are more prone to give unsafe responses when prompted in non-English languages. This work represents a significant step toward closing the safety gap in multilingual LLMs by enabling the development of culturally aware safety guard models.
@@ 1 Introduction
Figure 3: Overview of the Nemotron-Content-Safety-Dataset-Multilingual-v1 (CultureGuard dataset), showing sample counts across train, test, and validation splits, derived from various sources through the CultureGuard pipeline and Jail-break Synthetic Data Generation (JB SDG).
@@
In response to these challenges, we present CultureGuard, a novel framework designed for curating culturally aligned, high-quality safety datasets across multiple languages. This method is purely based on synthetic data generation, leveraging the English generation capabilities of LLMs and strong translation methods. Importantly, it does not require human input, which presents a scalable approach for targeting multiple languages or regions. Our technical approach employs a four-stage pipeline, encompassing cultural data segregation, cultural data adaptation, machine translation, and quality filtering. Cultural adaptation of the samples using LLMs to the target culture is central to the CultureGuard approach; its core components are shown in Figure 2. It is important to note that since this adaptation is performed on the English source text prior to translation, our approach focuses on ideational and social cultural elements, while linguistic aspects are not explicitly covered Liu et al. (2025). We also present a novel cross-lingual consistency filter to remove low-quality samples post-translation. This pipeline facilitates the systematic conversion and expansion of the English Nemotron-Content-Safety-Dataset-V2 (formerly known as Aegis 2.0) safety dataset Ghosh et al. (2025) into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese (Mandarin). Additionally, we synthetically curate jail-breaking prompt-response pairs using LLMs to enhance the detection capabilities for such adversarial inputs. Safety labels for these synthetic pairs are computed via a jury of LLMs. The final dataset, termed Nemotron-Content-Safety-Dataset-Multilingual-v1111Nemotron-Content-Safety-Dataset-Multilingual-v1, consists of a total of 386,661 samples across 9 languages (including English) as shown in Figure 3. The dataset is based on the English Nemotron-Content-Safety-Dataset-V2 and uses the same safety risk taxonomy.
@@
This comprehensive dataset is then used to fine-tune the multilingual Llama-3.1-8B-Instruct model using a LoRA-based approach, resulting in the Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1222Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 model. The empirical evaluation of this model quantitatively demonstrates its efficacy, outperforming all publicly available LLMs on a diverse multilingual test dataset. Specifically, our fine-tuned model achieves a 31.06% relative improvement in multilingual scores compared to the publicly released Llama-Nemotron-Safety-Guard-V2 model (formerly known as Llama-3.1-AegisGuard) and approximately 1.28% over PolyGuard-Qwen, despite PolyGuard’s dataset being three times larger and non-commercially usable. Figure 1 illustrates this performance comparison on a per-language basis. This establishes Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 as the sole multilingual safety guard model currently available with a commercially-friendly license. This marks a significant stride towards bridging the safety gap in multilingual LLMs by enabling the development of truly culturally-aware safety guard models.
@@
  • Using the CultureGuard approach, we curate the Nemotron-Content-Safety-Dataset-Multilingual-v1 dataset with  386k samples in 9 languages following the Nemotron-Content-Safety-Dataset-V2 taxonomy. The taxonomy is structured into 12 top-level hazard categories along with 9 fine-grained subcategories (excluding the safe and needs caution categories).
  • This dataset is used to train Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1, which is a state-of-the-art model outperforming all other multilingual LLMs and Guard models on multiple multilingual benchmarks. The dataset and model are released publicly.
@@
For clarity and simplicity throughout this paper, we adopt the following terminology. The suite of models trained in this work is collectively referred to as the CultureGuard model variants. Among these, we identify Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 as the best-performing variant, which utilizes all the components proposed in this work. The dataset, Nemotron-Content-Safety-Dataset-Multilingual-v1, is referred to as the CultureGuard dataset. More broadly, the name CultureGuard also represents our proposed pipeline for curating such a multilingual cultural dataset.
@@ 3.1.4 Quality Filtering
To mitigate the second problem of translation mistakes changing meaning, we employ FAITH-based filtering. We use an LLM Judge to rate the translations based on five crucial aspects: Fluency, Accuracy, Idiomaticity, Terminology, and Handling of Format (FAITH). We use a low filtering threshold (3.5/5) here so that minimal samples are being filtered out.
@@
Table 1: Performance comparison of safety guard models on various multilingual safety benchmarks, measured by harmful-F1 score. CultureGuard model variations in this table refer to the Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 model. CultureGuard (vanilla) utilizes a translation and filtering pipeline. CultureGuard additionally incorporates cultural adaptation. JB SDG refers to the use of multilingual jailbreak Synthetic Data Generation, and QF denotes quality filtering. Benchmark abbreviations: CultureGuard (Nemotron-Content-Safety-Dataset-Multilingual-v1), PGPrompts (PolyGuardPrompts). Nemotron-Safety-Guard-V2 refers to Llama-Nemotron-Safety-Guard-V2.
@@ 3.3 Training and Evaluation
For evaluation, we utilize test sets that were curated as a part of this work, alongside several key multilingual safety benchmarks. The test sets curated as a part of Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 include:
@@ 4 Results and Discussion
  • CultureGuard + JB SDG + QF: This is our final model configuration, which further enhances the CultureGuard + JB SDG setup with Quality Filtering (QF). The model trained using this configuration is referred to as Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1.
In addition to these ablation studies, our evaluation is contextualized by a comparison against general-purpose LLMs serving as guard models, as presented in Table 9. Furthermore, we benchmark the inherent multilingual safety of leading open LLMs, with results detailed in Table 4.
@@ 4.5 Multilingual Safety Benchmarking of Open LLMs
These results collectively establish Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1 (CultureGuard + JB SDG + QF) as a state-of-the-art multilingual safety guard model, demonstrating superior performance across a wide array of content safety and adversarial benchmarks.
@@ 5 Conclusion
Leveraging the CultureGuard approach, we curated the Nemotron-Content-Safety-Dataset-Multilingual-v1 (386,661 samples across 9 languages) and used it to train Llama-3.1-Nemotron-Safety-Guard-Multilingual-8B-v1. This model consistently outperforms major publicly available LLMs on diverse multilingual safety benchmarks, demonstrating a 31.06% improvement over English Llama-Nemotron-Safety-Guard-V2 and 1.28% over multilingual PolyGuard-Qwen, despite the latter’s substantially larger dataset and non-commercial usability. Notably, our model is currently the sole commercially-friendly multilingual safety guard model available.
@@ Acknowledgements
This work would not have been possible without contributions from many people at NVIDIA. To mention a few: Pinky Xu, Jayda Ritchie, Monika Katariya, Suchismita Sahu, Suseella Panguluri, and Christopher Parisien.

### v3

@@ Abstract
The increasing use of Large Language Models (LLMs) in agentic applications highlights the need for robust safety guard models. While content safety in English is well-studied, non-English languages lack similar advancements due to the high cost of collecting culturally aligned labeled datasets. We present CultureGuard, a novel solution for curating culturally aligned, high-quality safety datasets across multiple languages. Our approach introduces a four-stage synthetic data generation and filtering pipeline: cultural data segregation, cultural data adaptation, machine translation, and quality filtering. This pipeline enables the conversion and expansion of the Nemotron-Content-Safety-Dataset-V2 English safety dataset into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese. The resulting dataset, Nemotron-Safety-Guard-Dataset-v3, comprises 386,661 samples in 9 languages and facilitates the training of Llama-3.1-Nemotron-Safety-Guard-8B-v3 via LoRA-based fine-tuning. The final model achieves state-of-the-art performance on several multilingual content safety benchmarks. Furthermore, we show our moderately multilingual fine-tuning enables robust cross-lingual transfer and strong zero-shot generalization to unseen languages. We also benchmark the latest open LLMs on multilingual safety and observe that these LLMs are more prone to give unsafe responses when prompted in non-English languages. This work advances multilingual LLM safety by enabling the development of culturally aware safety guard models.
@@ 1 Introduction
Figure 3: Overview of the Nemotron-Safety-Guard-Dataset-v3 (CultureGuard dataset), showing sample counts across train, test, and validation splits, derived from various sources through the CultureGuard pipeline and Jail-break Synthetic Data Generation (JB SDG).
@@
In response to these challenges, we present CultureGuard, a novel framework designed for curating culturally aligned, high-quality safety datasets across multiple languages. This method is purely based on synthetic data generation, leveraging the English generation capabilities of LLMs and strong translation methods. Importantly, it does not require human input, which presents a scalable approach for targeting multiple languages or regions. Our technical approach employs a four-stage pipeline, encompassing cultural data segregation, cultural data adaptation, machine translation, and quality filtering. Cultural adaptation of the samples using LLMs to the target culture is central to the CultureGuard approach; its core components are shown in Figure 2. It is important to note that since this adaptation is performed on the English source text prior to translation, our approach focuses on ideational and social cultural elements, while linguistic aspects are not explicitly covered Liu et al. (2025). We also present a novel cross-lingual consistency filter to remove low-quality samples post-translation. This pipeline facilitates the systematic conversion and expansion of the English Nemotron-Content-Safety-Dataset-V2 (formerly known as Aegis 2.0) safety dataset Ghosh et al. (2025) into eight distinct languages: Arabic, German, Spanish, French, Hindi, Japanese, Thai, and Chinese (Mandarin). Additionally, we synthetically curate jail-breaking prompt-response pairs using LLMs to enhance the detection capabilities for such adversarial inputs. Safety labels for these synthetic pairs are computed via a jury of LLMs. The final dataset, termed Nemotron-Safety-Guard-Dataset-v3111Nemotron-Safety-Guard-Dataset-v3, consists of a total of 386,661 samples across 9 languages (including English) as shown in Figure 3. The dataset is based on the English Nemotron-Content-Safety-Dataset-V2 and uses the same safety risk taxonomy.
@@
This comprehensive dataset is then used to fine-tune the multilingual Llama-3.1-8B-Instruct model using a LoRA-based approach, resulting in the Llama-3.1-Nemotron-Safety-Guard-8B-v3222Llama-3.1-Nemotron-Safety-Guard-8B-v3 model. The empirical evaluation of this model quantitatively demonstrates its efficacy, outperforming all publicly available LLMs on a diverse multilingual test dataset. Specifically, our fine-tuned model achieves a 31.06% relative improvement in multilingual scores compared to the publicly released Llama-Nemotron-Safety-Guard-V2 model (formerly known as Llama-3.1-AegisGuard) and approximately 1.28% over PolyGuard-Qwen, despite PolyGuard’s dataset being three times larger and non-commercially usable. Figure 1 illustrates this performance comparison on a per-language basis. This establishes Llama-3.1-Nemotron-Safety-Guard-8B-v3 as the sole multilingual safety guard model currently available with a commercially-friendly license. Moreover, our evaluation also extends to languages unseen during fine-tuning, where the model maintains comparable performance. This demonstrates strong zero-shot generalization, expanding its effective support to over 20 languages. This marks a significant stride towards bridging the safety gap in multilingual LLMs by enabling the development of truly culturally-aware safety guard models.
@@
  • Using the CultureGuard approach, we curate the Nemotron-Safety-Guard-Dataset-v3 dataset with  386k samples in 9 languages following the Nemotron-Content-Safety-Dataset-V2 taxonomy. The taxonomy is structured into 12 top-level hazard categories along with 9 fine-grained subcategories (excluding the safe and needs caution categories).
  • This dataset is used to train Llama-3.1-Nemotron-Safety-Guard-8B-v3, which is a state-of-the-art model outperforming all other multilingual LLMs and Guard models on multiple multilingual benchmarks. The dataset and model are released publicly.
  • We present an ablation study demonstrating that fine-tuning on a diverse set of languages (four or more) significantly enhances the model’s zero-shot generalization capabilities to languages unseen during training.
@@
For clarity and simplicity throughout this paper, we adopt the following terminology. The suite of models trained in this work is collectively referred to as the CultureGuard model variants. Among these, we identify Llama-3.1-Nemotron-Safety-Guard-8B-v3 as the best-performing variant, which utilizes all the components proposed in this work. The dataset, Nemotron-Safety-Guard-Dataset-v3, is referred to as the CultureGuard dataset. More broadly, the name CultureGuard also represents our proposed pipeline for curating such a multilingual cultural dataset.
@@ 3.1.4 Quality Filtering
To mitigate the second problem of translation mistakes changing meaning, we employ FAITH-based filtering Paul et al. (2025). We use an LLM Judge to rate the translations based on five crucial aspects: Fluency, Accuracy, Idiomaticity, Terminology, and Handling of Format (FAITH). We use a low filtering threshold (3.5/5) here so that minimal samples are being filtered out.
@@
Table 1: Performance comparison of safety guard models on various multilingual safety benchmarks, measured by harmful-F1 score. CultureGuard model variations in this table refer to the Llama-3.1-Nemotron-Safety-Guard-8B-v3 model. CultureGuard (vanilla) utilizes a translation and filtering pipeline. CultureGuard additionally incorporates cultural adaptation. JB SDG refers to the use of multilingual jailbreak Synthetic Data Generation, and QF denotes quality filtering. Benchmark abbreviations: CultureGuard (Nemotron-Safety-Guard-Dataset-v3), PGPrompts (PolyGuardPrompts). Nemotron-Safety-Guard-V2 refers to Llama-Nemotron-Safety-Guard-V2.
@@
# Training CultureGuard CultureGuard PGPrompts PGPrompts Languages (Prompt) (Response) (Prompt) (Response) 2 79.13 80.40 70.71 68.10 4 84.30 84.20 78.59 73.33 6 84.79 83.89 78.66 71.56 9 84.95 85.25 79.00 73.12 12 85.34 84.68 79.04 73.16
Table 3: Ablation study on the impact of multilingual fine-tuning on zero-shot generalization performance. We report the average harmful F1-score on a set of 17 languages [^8] The 17 test languages differ slightly between datasets. Overall 20 unique languages are tested.
CultureGuard: en, ar, de, es, fr, hi, ja, th, zh, it, ko, nl, cs, da, fi, iw, pt-BR; PGPrompts: en, ar, de, es, fr, hi, ja, th, zh, it, ko, nl, cs, pl, pt, ru, sv
The training language sets were expanded as follows:
2: en, hi
4: en, ar, hi, ja
6: en, ar, es, hi, ja, zh
9: en, ar, de, es, fr, hi, ja, th, zh
12: en, ar, de, es, fr, hi, ja, th, zh, it, ko, nl
; all other analyses in this work focus on a 9-language set.
@@ 3.3 Training and Evaluation
For evaluation, we utilize test sets that were curated as a part of this work, alongside several key multilingual safety benchmarks. The test sets curated as a part of Llama-3.1-Nemotron-Safety-Guard-8B-v3 include:
@@ 4 Results and Discussion
  • CultureGuard + JB SDG + QF: This is our final model configuration, which further enhances the CultureGuard + JB SDG setup with Quality Filtering (QF). The model trained using this configuration is referred to as Llama-3.1-Nemotron-Safety-Guard-8B-v3.
In addition to these ablation studies, the zero-shot generalization results are presented in Table 8. Our evaluation is contextualized by a comparison against general-purpose LLMs serving as guard models, as presented in Table 11. Furthermore, we benchmark the inherent multilingual safety of leading open LLMs, with results detailed in Table 5.
@@ 4.3 Impact of Jail-Breaking Synthetic Data and Filtering
4.4 Zero-Shot Generalization to Unseen Languages
To evaluate zero-shot generalization, we conducted an ablation study by training models on language sets of increasing size (from 2 to 12) using the CultureGuard pipeline and testing them on a 20-language suite. As detailed in Table 8, the results demonstrate robust zero-shot capabilities; performance on unseen languages was comparable to that on the seen languages (Table 1). We observed a significant performance gain when expanding from two to four training languages, followed by diminishing returns. This indicates that a moderately diverse training set is crucial for inducing robust cross-lingual transfer of safety knowledge. Based on this performance trade-off, our 9-language model was selected as the optimal configuration, enabling it to effectively support over 20 languages through strong generalization. Further analysis across a broader set of unseen languages is left as an area for future work.
@@ 4.6 Multilingual Safety Benchmarking of Open LLMs
These results collectively establish Llama-3.1-Nemotron-Safety-Guard-8B-v3 (CultureGuard + JB SDG + QF) as a state-of-the-art multilingual safety guard model, demonstrating superior performance across a wide array of content safety and adversarial benchmarks.
@@ 5 Conclusion
Leveraging the CultureGuard approach, we curated the Nemotron-Safety-Guard-Dataset-v3 (386,661 samples across 9 languages) and used it to train Llama-3.1-Nemotron-Safety-Guard-8B-v3. This model consistently outperforms major publicly available LLMs on diverse multilingual safety benchmarks, demonstrating a 31.06% improvement over English Llama-Nemotron-Safety-Guard-V2 and 1.28% over multilingual PolyGuard-Qwen, despite the latter’s substantially larger dataset and non-commercial usability. Beyond these benchmarks, the model exhibits strong zero-shot generalization, enabling it to effectively support over 20 languages despite being trained on only nine. Notably, our model is currently the sole commercially-friendly multilingual safety guard model available.
@@ Acknowledgements
This work would not have been possible without contributions from many people at NVIDIA. To mention a few: Pinky Xu, Jayda Ritchie, Monika Katariya, Suchismita Sahu, Suseella Panguluri, Shyamala Prayaga, Isabel Hulseman, and Christopher Parisien.
@@ Appendix A Appendix
Models Average en ar th zh Nemotron-Safety-Guard-V2 59.14 91.81 43.49 46.35 54.92 Llama-Guard-3-8B 64.55 68.68 62.54 66.03 60.95 Llama-Guard-4-12B 58.71 64.06 50.48 59.37 60.95 Granite Guardian 3.1 8B 79.42 96.09 73.33 62.86 85.40 Granite Guardian 3.3 8B (reasoning off) 71.95 95.73 67.30 39.37 85.40 Granite Guardian 3.3 8B (reasoning on) 77.41 95.37 73.02 56.19 85.08 PolyGuard-Qwen 89.58 96.09 84.76 86.98 90.48 CultureGuard (vanilla) 87.62 91.10 86.03 86.98 86.35 CultureGuard 89.02 93.24 86.35 88.57 87.94 CultureGuard + JB SDG 89.72 95.37 85.71 88.57 89.21 CultureGuard + JB SDG + QF 92.28 96.09 89.21 91.11 92.70
Table 14: Harmful content classification performance (harmful-recall) on MultiJail test set (prompt classification).
@@
Models Average en ar de fr hi ja zh Nemotron-Safety-Guard-V2 24.53 70.11 14.39 18.29 20.07 13.21 11.21 24.39 Llama-Guard-3-8B 26.83 47.61 22.68 22.79 24.21 22.25 23.29 25.00 Llama-Guard-4-12B 25.15 41.28 25.50 23.18 23.64 18.58 22.79 21.07 Granite Guardian 3.1 8B 51.13 77.59 42.18 43.04 46.82 51.67 47.32 49.32 Granite Guardian 3.3 8B (reasoning off) 42.99 75.00 31.71 39.50 42.43 32.08 37.43 42.75 Granite Guardian 3.3 8B (reasoning on) 44.95 74.81 37.04 40.25 42.64 36.75 38.93 44.21 PolyGuard-Qwen 50.34 80.36 45.18 44.43 45.54 43.67 47.14 46.04 CultureGuard (vanilla) 44.57 70.79 41.14 38.96 40.25 40.08 40.25 40.50 CultureGuard 47.99 73.66 43.75 41.36 42.00 45.67 45.43 44.04 CultureGuard + JB SDG 48.33 77.30 43.04 42.46 43.21 42.46 44.43 45.43 CultureGuard + JB SDG + QF 53.27 80.65 49.75 47.32 46.71 48.96 49.96 49.50
Table 16: Harmful content classification performance (harmful-recall) on XSafety test set (prompt classification).
@@
Models Average en ar es fr hi Nemotron-Safety-Guard-V2 56.98 81.46 27.89 63.30 57.20 55.08 Llama-Guard-3-8B 68.20 62.51 57.67 72.51 71.83 76.50 Llama-Guard-4-12B 54.05 46.10 49.22 57.93 56.46 60.55 Granite Guardian 3.1 8B 88.45 91.79 79.56 93.09 91.88 85.90 Granite Guardian 3.3 8B (reasoning off) 84.30 92.40 65.00 92.97 92.00 79.13 Granite Guardian 3.3 8B (reasoning on) 85.94 92.91 69.22 91.43 92.87 83.28 PolyGuard-Qwen 91.35 90.78 87.67 95.91 92.99 89.40 CultureGuard (vanilla) 88.44 81.16 85.78 92.97 89.42 92.90 CultureGuard 90.41 84.90 85.33 95.91 91.27 94.64 CultureGuard + JB SDG 91.47 87.64 88.56 95.40 91.88 93.88 CultureGuard + JB SDG + QF 93.81 90.17 93.67 96.55 93.48 95.19
Table 20: Harmful content classification performance (harmful-recall) on Aya Red-teaming test set (prompt classification).